OpenStudy (anonymous):

A railroad tunnel is shaped like a semiellipse as shown below. A semiellipse is shown on the coordinate plane with vertices on the x axis and one point of intersection with the positive y axis. The height of the tunnel at the center is 58 ft and the vertical clearance must be 29 ft at a point 21 ft from the center. Find an equation for the ellipse

OpenStudy (amistre64):

what is the standard form of an ellipse equation, itll give us something to start filling in

OpenStudy (anonymous):

(x-h^2)/a^2+(y-k)^2/b^2=1

OpenStudy (amistre64):

good, and the h,k parts are 0 since we want to center it. the under y part is the height at the center, squared.

OpenStudy (amistre64):

all thats left to do is determine the value of a. any ideas?

OpenStudy (anonymous):

a is the y-value?

OpenStudy (amistre64):

b is the height stated in the problem, it is the distance from center to the y vertex ...

OpenStudy (amistre64):

when y=b, then b^2/b^2 = 1

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so, h=k=0 to center it at the origin of the graph ... and b=58 as stated. any ideas to determine a?

OpenStudy (amistre64):

we have reduced it to one equation with 1 unknown, and a known point.

OpenStudy (amistre64):

use the point and solve for a

OpenStudy (anonymous):

what is the reduced equation again?

OpenStudy (amistre64):

i hve no mouse to copy and paste with. rewrite your standard equation but with h=k=0 and b=58

OpenStudy (anonymous):

(x-h)^2/a^2+(y-k)^2/b^2=1 (x-0)^/a^2+(y-0)^2/58^2 x^2/a^2+y^2/3364=1

OpenStudy (amistre64):

good, now we need the point (x=21,y=29) if memory serves, so input those and solve for a

OpenStudy (anonymous):

(21)^2/a^2+(29)^2/3364=1 441/a^2+841/3364=1

OpenStudy (amistre64):

so far so good, the rest is basic algebra ... subtract, multiply, and divide as needed

OpenStudy (amistre64):

y reduces to 1/4 to help out lol

OpenStudy (amistre64):

the y parts reduce to 1/4 that is .... i get lazy typing at times

OpenStudy (anonymous):

441/a^2+1/4=1 a^2=558 a=24.248

OpenStudy (amistre64):

14 sqrt3 but yeah

OpenStudy (amistre64):

it should suffice for a^2 = 588 since a^2 is part of the formula

OpenStudy (anonymous):

So, the equation is:\[x^2/14\sqrt3+1/4=1\]

OpenStudy (amistre64):

no, x^2/a^2 + y^2/b^2 = 1 is the formula. we simply used it to find a^2, and b^2 is given as 58^2

OpenStudy (anonymous):

Ok, so it is x^2/14sqrt3+y^2/58^2=1?

OpenStudy (amistre64):

a^2 = 588 which is what the formula requires. finding a was a pointless endeavor

OpenStudy (amistre64):

x^2/5988 + y^2/58^2 = 1 is what we are looking for to satisfy the given information

OpenStudy (amistre64):

pfft, me fat finngers ... x^2/588 + ...

OpenStudy (anonymous):

Ok,Thank You